2 charges q and -2q are separated by a distance d.If the electric intensity at the site of q is E then electric field at the site of -2q is
1)E
2)E/2
3)-2E
4)-E/2

Grade:11

2 Answers

Piyush Kumar Behera
4 years ago
the answer of the above question is attached below in the image.The correct option is 4) -E/2
 
Piyush Kumar Behera
IIT Dhanbad
Khimraj
4 years ago
electric intensity at the site of q = K*(-2q)/d2 = E
then electric intensity at the site of -2q = K*q/d2 = -E/2
Hope it clears.................